# DreamMotion: Space-Time Self-Similarity Score Distillation for Zero-Shot Video Editing

### v1

@@ 4.2.1 Setup
In this experiment, we utilize the 8-frame videos from the previously assembled text-video pairs. Additionally, we benefit from Show-1 [63], an open-source, cascaded video diffusion model. As detailed in Sec. 3.5, we compose our cascaded pipeline comprising Keyframe Generation, Temporal Interpolation, and Spatial Super Resolution, with all modules operating in pixel-space. Our method is implemented during the initial keyframe generation stage. During keyframe optimization, these input videos undergo resizing to a resolution of 80x128 pixels, with the optimization process taking approximately 3 minutes on a single A100 GPU. Following the optimization, the frame interpolation and super-resolution modules expand the output keyframes temporally and spatially, respectively.
@@ 4.2.2 Baselines
To our knowledge, VMC [20] stands out as the sole video editing approach utilizing a cascaded video diffusion pipeline. VMC adapts temporal attention layers within the keyframe generation module, leveraging their novel motion distillation objective. For comparison purpose, we introduce an additional variant that employs direct inference using the cascaded pipeline with modified target text, starting from the DDIM inverted latents [47].
@@ 4.2.3 Qualitative Results
We qualitatively compare our method against the baselines in Fig. 8. Our method generate videos that match the structure and layout of the input video while faithfully adhering to the edit prompt, while other methods struggle to maintain the structural and motion integrity of the original video. As all three methods employ unaltered temporal interpolation and super-resolution models following the generation of keyframes, they commonly produce temporally consistent videos. Refer to Supplementary Material and project page for full results.
@@ 4.2.4 Quantitative Results
Adopting the metrics outlined in Sec. 4.1, we conduct a quantitative comparison of our method against baseline approaches, detailed in Tab. 1. Notably, our approach demonstrated substantial superiority in Structure and Motion Preservation (SM-Preserve).
@@ 4.3 Ablation Studies
In Fig. 6, we evaluate the impact of using binary masks to selectively filter gradients during $\mathcal{L}_{\text{V-DDS}}$ update. The results demonstrate that filtering gradients responsible for appearance injection enhances the precision of video editing and improves visual fidelity, while avoiding issues of blurriness and saturation.
We next ablate the necessity of our self-similarity guidance. Fig. 3 illustrates the optimization progress with and without our self-similarity alignments. The process begins with the initial input video (top row). Solely using $\mathcal{L}_{\text{V-DDS}}$ for appearance injection (left) leads to the accumulation of structural errors as optimization progresses, resulting in motion deviation in the final output. However, when the process is regularized by the spatial and temporal self-similarities (right), edited videos maintain the structure and motion fidelity throughout the optimization. Additionally, in Fig. 9, we illustrate video editing results under different optimization setups: (i) $\mathcal{L}_{\text{V-DDS}}$. (ii) $\mathcal{L}_{\text{V-DDS}}+\mathcal{L}_{\text{S-SSM}}$. (ii) $\mathcal{L}_{\text{V-DDS}}+\mathcal{L}_{\text{S-SSM}}+\mathcal{L}_{\text{T-SSM}}$. Absence of spatial self-similarity loss leads to inconsistency in object structures across frames. For instance, the shape of a bird’s wing varies, creating visible discrepancies, as shown in Fig. 9-left. While aligning spatial self-similarity with the original video preserves structural integrity, it may generate artifacts in optimized areas. However, these artifacts are efficiently addressed through the addition of temporal self-similarity guidance.
Figure 10: Limitation. Our method limits its ability to produce videos that necessitate substantial structural alterations.
@@ 5 Conclusion
In this work, we have addressed the intricate challenge of diffusion-based video editing, a domain where formulating temporally consistent, real-world motion remains a notable obstacle. DreamMotion introduced score distillation based optimization to text-to-video diffusion models, marking a departure from traditional, ancestral sampling based video editing. Our framework adeptly incorporated new content as specified by target text descriptions using Video Delta Denoising Score, while preserving the the structural integrity and motion of the original video via a novel space-time self-similarity alignment. Through rigorous validation in both cascaded and non-cascaded video diffusion settings, our approach has proven superior in maintaining the essence of the original video while seamlessly integrating desired alterations. Regarding limitations, our framework is designed to preserve structural integrity of the original video, and as such, it is not suited for edits that require significant structural changes (see Fig. 10).
Ethics Statement. Our work is based on generative models that carry the risk of being repurposed for unethical uses, including the creation of misleading or counterfeit content, which may yield adverse societal impacts.
@@ 0.A.1 Video Editing using Diffusion Models
Creating videos from textual descriptions necessitates ensuring realistic and temporally consistent motion, posing a unique set of challenges compared to the text-driven image generative scenarios. Before the advent of publicly accessible text-to-video diffusion models, Tune-A-Video [60] was at the forefront of one-shot based video editing. They proposed to inflate image diffusion model to pseudo video diffusion model by appending temporal modules to image diffusion model [40] and reformulating spatial self-attention into spatio-temporal self-attention, facilitating inter-frame interactions. However, the inflation often falls short in achieving consistent and complete motion, as motion preservation relies implicitly on the attention mechanism during inference. Thus, the attention projection matrices within U-Net are often fine-tuned on the input videos [60, 28, 67].
Utilizing explicit visual signals to steer the video denoising process is another common technique. Pix2Video [4], for instance, incorporates self-attention features from previously edited frames into the self-attention layer of the current frame, whereas FateZero[37] retains all intermediate attention maps obtained during the input video inversion phase and integrates them into the attention layers throughout the editing stage. stores all intermediate attention maps derived during the input video inversion and injects them to the attention layers during the editing phase. Others leverage pretrained image adapter networks for structurally consistent video generation. A notable example is ControlNet [64], which has been modified to accommodate a series of explicit structural indicators such as depth and edge maps. Ground-A-Video [21] takes this a step further by adapting both ControlNet and GLIGEN [26] for video editing, utilizing spatially-continuous depth maps and spatially-discrete bounding boxes.
Despite the availability of open-source text-to-video (T2V) diffusion models [55, 50, 5, 57, 63], recent endeavors frequently adopt a self-supervised strategy of fine-tuning pretrained video generative models on an input video, to accurately capture intricate, real-world motion. More specifically, several studies attempt to disentangle appearance and motion elements of videos during the self-supervised fine-tuning. For example, [68, 59, 66] split the fine-tuning phase into two distinct pathways: one dedicated to integrating the subject’s appearance into spatial modules, and the other aimed at embedding motion dynamics of a video into temporal modules within the T2V model. Additionally, other studies [20, 29] attempt to extract and learn motion information from a single or a few reference videos. VMC [20], for instance, propose to distill the motion within a video by calculating the residual vectors between consecutive frames, and refine temporal attention layers in cascaded video diffusion models.
@@ 0.A.2 Visual Generation using Score Distillation Sampling
Score Distillation Sampling (SDS) [36], also known as Score Jacobian Chaining, has become the go-to method for text-to-3D generation in recent years [58, 30, 27, 18, 52, 44, 6, 33]. DreamFusion [36] first proposed to distill the generative prior of pretrained text-to-image models and optimize a parametric image synthesis model, such as NeRF [31]. Despite its success, SDS often produces images that are overly saturated, blurry, and lack detail, largely due to the use of high CFG values [58].
To address these challenges, a range of derivative methods have been proposed [58, 30, 27, 18, 11, 32]. Specifically, in the context of accurate image editing, DDS [11] incorporates additional reference branch with corresponding text to refine the noisy gradient of SDS. Hifa [69], instead, utilizes an estimated clean image rather than the predicted noise to compute denoising scores. In our work, we employ a straightforward yet effective mask condition to refine DDS-generated gradients, allowing us to inject particular appearance to the video. We further ensures the preservation of the video’s original structure and motion through the novel regularization of space-time self-similarity alignment.
@@ Appendix 0.C User Study Interface
We carried out human evaluations to assess various methods based on three key apsects: Edit Accuracy, Frame Consistency, and Sturcture & Motion Preservation. Initially, we present the input video alongside its text description, as shown in Figure 11. Subsequently, we display target text with anonymized videos generated by each method and ask participants to evaluate them across the aforementioned three criteria. The human evaluation results, detailed in Table 1 of the manuscript, unequivocally highlight superiority of DreamMotion in both video diffusion frameworks.
Appendix 0.D Additional Results
This section is dedicated to presenting additional outcomes of DreamMotion. Figure 12 offers a comprehensive view of the results from Figure 6 in the main paper, demonstrating the effect of masking DDS-driven gradients. Annotations within the input video frames indicate the masks used. In Figure 13, we present the progress of DreamMotion optimization by visualizing intermediate output videos. Figures 14, 15, and 16 showcase input and corresponding edited videos generated with DreamMotion on Zeroscope, using various target prompts. To accommodate space constraints, only odd or even frames from 16-frame videos are selected for display. Figures 17, 18, and 19 feature videos edited by DreamMotion on the Show-1 Cascade model [63], with the left columns displaying 8-frame input videos and the adjacent columns showing 29-frame output videos. Our qualitative results are uploaded on our project page.

### v2

@@ 4.2.1 Setup
In this experiment, we utilize the 8-frame videos from the previously assembled text-video pairs. Additionally, we benefit from Show-1 [66], an open-source, cascaded video diffusion model. As detailed in Sec. 3.5, we compose our cascaded pipeline comprising Keyframe Generation, Temporal Interpolation, and Spatial Super Resolution, with all modules operating in pixel space. Our method is implemented during the initial keyframe generation stage. During keyframe optimization, these input videos undergo resizing to a resolution of 80x128 pixels, with the optimization process taking approximately 3 minutes on a single A100 GPU. Following the optimization, the frame interpolation and super-resolution modules expand the output keyframes temporally and spatially, respectively.
@@ 4.2.2 Baselines
To our knowledge, VMC [20] stands out as the sole video editing approach utilizing a cascaded video diffusion pipeline. VMC adapts temporal attention layers within the keyframe generation module, leveraging their novel motion distillation objective. For comparison purposes, we introduce an additional variant that employs direct inference using the cascaded pipeline with modified target text, starting from the DDIM inverted latents [50].
@@ 4.2.3 Qualitative Results
We qualitatively compare our method against baselines in Fig. 8. DreamMotion generates videos that match the structure and layout of the input video while adhering to the edit prompt, while other methods struggle to maintain the structural and motion integrity of the original video. Since all three methods use unaltered temporal interpolation and super-resolution models after the generation of keyframes, they commonly produce temporally consistent videos. For comprehensive results, please refer to the appendix.
@@ 4.2.4 Quantitative Results
Adopting the metrics outlined in Sec. 4.1, we compare our method quantitatively against baseline approaches, detailed in Tab. 2. Notably, our approach demonstrated substantial superiority in Structure and Motion Preservation (SM-Preserve).
@@
Text-Align Frame-Con Motion-Fidelity Frame-LPIPS
Ours wo $\mathcal{L}_{\text{S-SSM}}+\mathcal{L}_{\text{T-SSM}}$ 0.8202 0.9648 0.8426 0.3247
Ours wo $\mathcal{L}_{\text{T-SSM}}$ 0.8114 0.9567 0.9011 0.3186
Ours wo masks 0.8180 0.9695 0.8653 0.3416
Ours (full) 0.8209 0.9726 0.9259 0.3042
Table 3: Quantitative ablation. We demonstrate the impact of each factor by removing individual losses and masking conditions.
@@ 4.3 Ablation Studies
In Fig. 6, we evaluate the impact of using bounding box-driven masks to selectively filter gradients during $\mathcal{L}_{\text{V-DDS}}$ update. The results demonstrate that filtering gradients responsible for appearance injection enhances the precision of video editing and improves visual fidelity while avoiding issues of blurriness and saturation.
We next ablate the necessity of our self-similarity guidances. Fig. 3 illustrates the optimization progress with and without our self-similarity alignments. The process begins with the initial input video (top row). Solely using $\mathcal{L}_{\text{V-DDS}}$ for appearance injection (left) leads to the accumulation of structural errors as optimization progresses, resulting in motion deviation in the final output. However, when the process is regularized by the spatial and temporal self-similarities (right), edited videos maintain the structure and motion fidelity throughout the optimization. Additionally, in Fig. 9, we illustrate video editing results under different optimization setups: (i) $\mathcal{L}_{\text{V-DDS}}$. (ii) $\mathcal{L}_{\text{V-DDS}}+\mathcal{L}_{\text{S-SSM}}$. (ii) $\mathcal{L}_{\text{V-DDS}}+\mathcal{L}_{\text{S-SSM}}+\mathcal{L}_{\text{T-SSM}}$. The absence of spatial self-similarity loss leads to inconsistency in object structures across frames. For instance, the shape of a bird’s wing varies, creating visible discrepancies, as shown in Fig. 9-left. While aligning spatial self-similarity with the original video preserves structural integrity, it may generate artifacts in optimized areas. However, these artifacts are efficiently addressed through the addition of temporal self-similarity guidance. Lastly, Tab. 3 provides a quantitative analysis of each optimization term and masking condition.
Figure 10: Limitation. DreamMotion limits its ability to produce videos that necessitate substantial structural alterations.
@@ 5 Conclusion
In this work, we have addressed the intricate challenge of diffusion-based video editing, a domain where formulating temporally consistent, real-world motion remains a notable obstacle. DreamMotion introduced score distillation-based optimization to text-to-video diffusion models, marking a departure from traditional, ancestral sampling-based video editing. Our framework adeptly incorporated new content as specified by target text descriptions using the Video Delta Denoising Score, while preserving the the structural integrity and motion of the original video via a novel space-time self-similarity alignment. Through rigorous validation in both cascaded and non-cascaded video diffusion settings, our approach has proven superior in maintaining the essence of the original video while seamlessly integrating desired alterations. Regarding limitations, our framework is designed to preserve the structural integrity of the original video, and as such, it is not suited for edits that require significant structural changes (see Fig. 10).
Ethics Statement Our work is based on generative models that carry the risk of being repurposed for unethical uses, such as misleading content.
Acknowledgments
This work was supported by the National Research Foundation of Korea under Grant RS-2024-00336454. This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT, Ministry of Science and ICT) (No. 2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation). This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program (KAIST)). This research was supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2023. This research was supported by Field-oriented Technology Development Project for Customs Administration funded by the Korea government (the Ministry of Science & ICT and the Korea Customs Service) through the National Research Foundation (NRF) of Korea under Grant NRF2021M3I1A1097910.
@@ 0.A.1 Video Editing using Diffusion Models
Creating videos from textual descriptions necessitates ensuring realistic and temporally consistent motion, posing a unique set of challenges compared to text-driven image generative scenarios. Before the advent of publicly accessible text-to-video diffusion models, Tune-A-Video [63] was at the forefront of one-shot-based video editing. They proposed to inflate the image diffusion model to the pseudo video diffusion model by appending temporal modules to the image diffusion model [43] and reformulating spatial self-attention into spatio-temporal self-attention, facilitating inter-frame interactions. However, the inflation often falls short of achieving consistent and complete motion, as motion preservation relies implicitly on the attention mechanism during inference. Thus, the attention projection matrices within U-Net are often fine-tuned on the input videos [63, 29, 70]. Utilizing explicit visual signals to steer the video denoising process is another common technique. Pix2Video [4] and FateZero [40], for instance, inject intermediate attention maps during the editing phase, which are derived during the input video inversion. Others leverage pre-trained image adapter networks for structurally consistent video generation. A notable example is ControlNet [67], which has been modified to accommodate a series of explicit structural indicators such as depth and edge maps. Ground-A-Video [21] takes this a step further by adapting both ControlNet and GLIGEN [27] for video editing, utilizing spatially-continuous depth maps and spatially-discrete bounding boxes.
Despite the availability of open-source text-to-video (T2V) diffusion models [58, 53, 5, 60, 66], recent endeavors frequently adopt a self-supervised strategy of fine-tuning pre-trained video generative models on an input video, to accurately capture intricate, real-world motion. More specifically, several studies attempt to disentangle the appearance and motion elements of videos during the self-supervised fine-tuning. For example, [71, 62, 69] split the fine-tuning phase into two distinct pathways: one dedicated to integrating the subject’s appearance into spatial modules, and the other aimed at embedding motion dynamics of a video into temporal modules within the T2V model. Additionally, other studies [20, 31, 35] attempt to extract and learn motion information from a single or a few reference videos. VMC [20], for instance, proposes to distill the motion within a video by calculating the residual vectors between consecutive frames, and refine temporal attention layers in cascaded video diffusion models.
@@ 0.A.2 Visual Generation using Score Distillation Sampling
Score Distillation Sampling (SDS) [39], also known as Score Jacobian Chaining, has become the go-to method for text-to-3D generation in recent years [61, 32, 28, 18, 55, 47, 6, 36]. DreamFusion [39] first proposed to distill the generative prior of pre trained text-to-image models and optimize a parametric image synthesis model, such as NeRF [33]. Despite its success, SDS often produces images that are overly saturated, blurry, and lack detail, largely due to the use of high CFG values [61]. To address these challenges, a range of derivative methods have been proposed [61, 32, 28, 18, 11, 34]. Specifically, in the context of accurate image editing, DDS [11] incorporates an additional reference branch with corresponding text to refine the noisy gradient of SDS. Hifa [72], instead, utilizes an estimated clean image rather than the predicted noise to compute denoising scores. In our work, we employ a straightforward yet effective mask condition to refine DDS-generated gradients, allowing us to inject particular appearance into the video. We further ensures the preservation of the video’s original structure and motion through the novel regularization of space-time self-similarity alignment.
@@ Appendix 0.C User Study Interface
We carried out human evaluations to assess various methods based on three key aspects: Edit Accuracy, Frame Consistency, and Sturcture & Motion Preservation. Initially, we present the input video alongside its text description, as shown in Figure 11. Subsequently, we display target text with anonymized videos generated by each method and ask participants to evaluate them across the aforementioned three criteria. The human evaluation results, detailed in Table 1 of the manuscript, unequivocally highlight the superiority of DreamMotion in both video diffusion frameworks.
Figure 12: Additional qualitative comparison with DMT and Video-P2P.
Video-P2P DMT Ours (Zeroscope)
Motion-Fidelity 0.7384 0.8697 0.9259
Frame-LPIPS 0.3395 0.3078 0.3042
Table 4: Additional quantitative comparison with DMT and Video-P2P.
Appendix 0.D Additional Comparison
We additionally compared our method with two video editing techniques specifically designed for localized editing: Video-P2P [29] and Diffusion-Motion-Transfer (DMT) [65]. For qualitative comparison, see Fig. 12. For quantitative comparison in Tab. 4, we employed tracking-based motion fidelity score [65] and framewise LPIPS [29] to evaluate spatial consistency.
Appendix 0.E Additional Results
This section is dedicated to presenting additional outcomes of DreamMotion. Figure 13 offers a comprehensive view of the results from Figure 6 in the main paper, demonstrating the effect of masking DDS-driven gradients. Annotations within the input video frames indicate the masks used. In Figure 14, we present the progress of DreamMotion optimization by visualizing intermediate output videos. Figures 15, 16, and 17 showcase input and corresponding edited videos generated with DreamMotion on Zeroscope, using various target prompts. To accommodate space constraints, only odd or even frames from 16-frame videos are selected for display. Figures 18, 19, and 20 feature videos edited by DreamMotion on the Show-1 Cascade model [66], with the left columns displaying 8-frame input videos and the adjacent columns showing 29-frame output videos. Our qualitative results are uploaded on our project page.